The National Safety Council reported that 52% of American turnpike drivers are men. A sample of 300 cars traveling southbound on the New Jersey Turnpike yesterday revealed that 170 were driven by men. At the .01 significance level, can we conclude that a larger proportion of men were driving on the New Jersey Turnpike than the national statistics indicate? First state H0 & Ha
a) Is this a Z or t test
b) Test statistic =?
c) Critical value =?
d) P-value =?
e) Reject Ho: (yes or no)

Answer by stanbon(57387)

The National Safety Council reported that 52% of American turnpike drivers are men. A sample of 300 cars traveling southbound on the New Jersey Turnpike yesterday revealed that 170 were driven by men. At the .01 significance level, can we conclude that a larger proportion of men were driving on the New Jersey Turnpike than the national statistics indicate? First state H0 & Ha
a) Is this a Z or t test
Ans: Proportion tests are z tests
---------------------------------------
b) Test statistic = t(170/300) = (0.57-0.52)/sqrt[0.52*0.48/300] = 1.733
---------------------------------
c) Critical value for right tail test with alpha=1% = 2.339
------------------------------
d) P-value = P(t > 1.733) = tcdf(1.733,100,299) = 0.0421
--------------------------
e) Reject Ho: (yes or no)
Since the p-value is greater less than 1%, fail to reject Ho.
===================================
Cheers,
Stan H.
===================================